Mark has 7 more dimes than nickels. The value of his dimes is 10 cents more than 3 times the value of his nickels. How many of each coin does Mark have?
Answer by ikleyn(52776)

.
The value of nickels is 5*N  cents.

The value of dimes is 10*(N+7) cents.


The condition says:

    The value of his dimes = 3 times the value of his nickels + 10 cents,     or


10*(N+7) = 3*(5N) + 10,   or

10N + 70 = 15N + 10,   or

10N - 15N = 10 - 70,   or

-5N = - 60  ====>  N = 12.


Answer.  12 nickels and 12+7 = 19 dimes.

Solved.

There is entire bunch of lessons on coin problems
    - Coin problems
    - More Coin problems
    - Solving coin problems without using equations
    - Kevin and Randy Muise have a jar containing coins
    - Typical coin problems from the archive
    - More complicated coin problems
    - Solving coin problems mentally by grouping without using equations
    - Santa Claus helps solving coin problem
    - OVERVIEW of lessons on coin word problems
in this site.

You will find there the lessons for all levels - from introductory to advanced,
and for all methods used - from one equation to two equations and even without equations.

Read them and become an expert in solution of coin problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Coin problems".